1. Match the following aqueous solutions with the appropriate letter from the column on the right.

1.

0.18 m

MgI2

  

A.

Highest boiling point

2.

0.19 m

Fe(NO3)2

B.

Second highest boiling point

3.

0.16 m

CrCl2

C.

Third highest boiling point

4.

0.50 m

Sucrose(nonelectrolyte)

D.

Lowest boiling point

2. Match the following aqueous solutions with the appropriate letter from the column on the right.

1.

0.13 m

AlBr3

A.

Lowest freezing point

2.

0.24 m

CoSO4

B.

Second lowest freezing point

3.

0.17 m

Ba(NO3)2

C.

Third lowest freezing point

4.

0.56 m

Glucose(nonelectrolyte)

D.

Highest freezing point

3. Match the following aqueous solutions with the appropriate letter from the column on the right.

1.

0.18 m

MgI2

A.

Lowest freezing point

2.

0.19 m

Fe(NO3)2

B.

Second lowest freezing point

3.

0.16 m

CrCl2

C.

Third lowest freezing point

4.

0.50 m

Sucrose(nonelectrolyte)

D.

Highest freezing point
Answer #1

Answer

1)

1. 0.18m MgI2 : B second highest boiling point

2. 0.19m Fe(NO3)2 : A highest boiling point

3. 0.16m CrCl2 : D Lowest boiling point

4. 0.50m sucrose(nonelectrolyte) : C Third highest boiling point

2)

1. 0.13m AlBr3 : B second lowest freezing point

2. 0.24m CoSO4 : A highest freezing point

3.0.17m Ba(NO3)2 : C third lowest freezing point

4. 0.56m Glucose : D lowest freezing point

3)

1. 0.18m MgI2 : B Second lowest freezing point

2. 0.19m Fe(NO3)2 : A Lowest freezing point

3. 0.50m Sucrose (nonelectrolyte) : C Third lowest freezing point

4. 0.16m CrCl2 : D Highest freezing point

Explanation

i) Freezing point depression is calculated by the followning formula

∆Tf = Kf × b × i

where ,

Kf = freezing point depression constant which depends on solvent

b = molality of solute

i = van't Hoff factor

for same solvent , ∆Tf is depends on b×i

if b×i is higher freezing point depression would be higher and freezing point would be lower

For example consider question 2

i for AlBr3 = 4 ( because 4 ions after dissociation)

i for CoSO4 = 2 ( because 2 ions after dissociation )

i for Ba(NO3)2 = 3 ( because 3 ions after dissociation )

i for Glucose = 1 ( because no dissociation)

b × i for AlBr3 = 4× 0.13m = 0.52m

b× i for CoSO4 = 2 × 0.24m = 0.48m

b × i for Ba(NO3)2 = 3 × 0.17 = 0.51 m

b × i for Glucose = 1 × 0.56m = 0.56m

For glucose b× i value is high , so it would have highest freezing point depression and lowest freezing point

ii) Boiling point elevation is calculated by the following dormula

∆Tb = Kb × b × i

where,

Kb = boiling point elevation constant which depends on solvent

Therefore, for same solvent ∆Tb depends on b× i

If b× i is higher boiling point elevation will be higher and boiling point would higher